# ComplexTempQA: A Large-Scale Dataset for Complex Temporal Question Answering

### v1

@@ 1 Introduction
Temporal Question Answering (TQA) is about answering questions that require an understanding of and reasoning about temporal knowledge [17, 20]. This makes it substantially different from traditional Question Answering (QA) (see Figure 1). The complexity of questions users tend to ask to search systems nowadays has increased. Unlike traditional keyword-based searches, directly posed questions  [15] of increasing complexity are now essential for evaluating QA as well as TQA approaches, driven by the advancements in the current-generation LLMs.
The few existing datasets for TQA, such as HOTPOTQA [47], TORQUE [28], and TEQUILA [19] fall however short in several regards: First, these datasets are limited in size, typically containing only a few thousand questions. This poses significant challenges for effectively training LLMs, as illustrated in  [28, 18, 25, 29, 44, 27]. Second, while these datasets predominantly focus on specific question categories about entities or time periods, they lack comprehensive coverage across a wide variety of question types. Third, existing datasets generally include only straightforward questions and miss complex inquiries that necessitate unfolding and inference to generate accurate responses, thereby limiting trained models’ ability to handle sophisticated temporal reasoning tasks [28, 29, 44, 27]. Fourth, the datasets generally cover only short periods of time and lack features such as difficulty ratings for questions and the ability to filter data by specific time periods, which is essential for detailed and customized temporal studies.
@@ 2 Related Work
TORQUE [28] introduces questions emphasizing temporal relations such as "before," "after," and "start". The authors trained a model to evaluate questions specifically based on these temporal constraints. QA TempEval dataset [25] has been designed with the focus on temporal entities and temporal relations, which are easier to generate automatically. TEQUILA [19] uses temporal expressions like dates or implicit temporal signals such as "before" or "after." In addition to these works, TempQuestions [20] has been released as a temporal questions benchmark, containing 1,271 questions that are all temporal in nature, paired with their answers. That work provides a simple definition for temporal questions and demonstrated the need for further research on complex questions.
@@ 6 Evaluation
RAG QA.
Retrieval-augmented generation (RAG) [24, 2] combines the strengths of pre-trained language models and information retrieval systems to generate responses in question-answering settings. RAG retrieves relevant documents from a large corpus when a question is posed, providing these documents as additional context to a language model to generate a response based on both the question and the retrieved documents.
@@ 8 Conclusions
In this paper, we presented ComplexTempQA, a large dataset with over 100 million temporal question-answer pairs. ComplexTempQA significantly outperforms existing benchmarks in scope and coverage. Built on Wikipedia and Wikidata, the dataset covers questions from over three decades and offers a unique range of topics. We further introduced a taxonomy that categorizes the questions into attributes, comparisons, and count questions, with each question focusing on events, entities, and time periods. Another feature of ComplexTempQA is the high complexity of the questions, which requires skills such as across-temporal comparisons, temporal aggregation and multi-hop inference with temporal event ordering and entity recognition. Each question is accompanied by detailed metadata, including specific temporal domains, which enabled comprehensive evaluation and improvement of the temporal reasoning capabilities of large language models (LLMs).
Additionally, we evaluated various LLMs on our dataset, obtaining insights into their performance on temporal questions. Overall, ComplexTempQA serves as a valuable resource for question answering, information retrieval, and language understanding. It is also useful, for example, for testing LLMs’ temporal factual knowledge, identifying the hallucinations they produce, and assessing their reasoning ability over time.

### v2

@@ 1 Introduction
Temporal Question Answering (TQA) is about answering questions that require an understanding of and reasoning about temporal knowledge [18, 22, 21, 30]. This makes it substantially different from traditional Question Answering (QA) (see Figure 1). The complexity of questions users tend to ask to search systems nowadays has increased. Unlike traditional keyword-based searches, directly posed questions  [16] of increasing complexity are now essential for evaluating QA as well as TQA approaches, driven by the advancements in the current-generation LLMs.
The few existing datasets for TQA, such as TORQUE [30], ArchivalQA [47], ChroniclingAmericaQA [33] and TEQUILA [21] fall however short in several regards: First, these datasets are limited in size, typically containing only a few thousand questions. This poses significant challenges for effectively training LLMs, as illustrated in  [30, 20, 27, 31, 49, 29]. Second, while these datasets predominantly focus on specific question categories about entities or time periods, they lack comprehensive coverage across a wide variety of question types. Third, existing datasets generally include only straightforward questions and miss complex inquiries that necessitate unfolding and inference to generate accurate responses, thereby limiting trained models’ ability to handle sophisticated temporal reasoning tasks [30, 31, 49, 29]. Fourth, the datasets generally cover only short periods of time and lack features such as difficulty ratings for questions and the ability to filter data by specific time periods, which is essential for detailed and customized temporal studies.
@@ 2 Related Work
TORQUE [30] introduces questions emphasizing temporal relations such as "before," "after," and "start". The authors trained a model to evaluate questions specifically based on these temporal constraints. QA TempEval dataset [27] has been designed with the focus on temporal entities and temporal relations, which are easier to generate automatically. TEQUILA [21] uses temporal expressions like dates or implicit temporal signals such as "before" or "after." In addition to these works, TempQuestions [22] has been released as a temporal questions benchmark, containing 1,271 questions that are all temporal in nature, paired with their answers. That work provides a simple definition for temporal questions and demonstrated the need for further research on complex questions. To improve question answering for event-focused questions in temporal collections of news articles, Wang et al. [46] propose a large-scale question-answering model designed specifically for long-term news article collections. The authors focus mainly on introducing an effective approach for re-ranking articles by using document publication dates and temporal expressions embedded in text. Stricker [38] leverage a popular approach used for general question answering and answer extraction to find answers to temporal questions within paragraphs.
@@ 6 Evaluation
RAG QA.  Retrieval-augmented generation (RAG) [26, 2] combines the strengths of pre-trained language models and information retrieval systems to generate responses in question-answering settings. RAG retrieves relevant documents from a large corpus when a question is posed, providing these documents as additional context to a language model to generate a response based on both the question and the retrieved documents.
@@ 8 Conclusions
In this paper, we presented ComplexTempQA, a large dataset with over 100 million temporal question-answer pairs. ComplexTempQA significantly outperforms existing benchmarks in scope and coverage. Built on Wikipedia and Wikidata, the dataset covers questions from over three decades and offers a unique range of topics. We further introduced a novel taxonomy that categorizes the questions into attributes, comparisons, and count questions, with each question focusing on events, entities, and time periods. Another feature of ComplexTempQA is the high complexity of the questions, which requires skills such as across-temporal comparisons, temporal aggregation and multi-hop inference with temporal event ordering and entity recognition. Each question is accompanied by detailed metadata, including specific temporal domains, which enabled comprehensive evaluation and improvement of the temporal reasoning capabilities of large language models (LLMs).
Additionally, we evaluated various LLMs on our dataset, obtaining insights into their performance on temporal questions. Overall, ComplexTempQA serves as a valuable resource for developing sophisticated AI models and for advancing research in question answering, information retrieval, and language understanding. It is also useful, for example, for testing LLMs’ temporal factual knowledge, identifying the hallucinations they produce, and assessing their reasoning ability over time.
In the future we want to focus on extending datasets to other time periods and other cmplex question types, as well as consider insights from collective memory related studies [13, 19] to better select events and time periods for question generation.